OpenStudy (anonymous):

Simplify: 3 square root of 10 end root plus 7 square root of 15 end root minus 6 square root of 10 end root minus 4 square root of 15 3 square root of 10 end root minus 3 square root of 15 3 square root of 15 end root minus 3 square root of 10 3 square root of 30 end root minus 3 square root of 20 3 square root of 20 end root minus 3 square root of 30

OpenStudy (anonymous):

will give medal to who ever solves it

OpenStudy (anonymous):

@RyGuy

OpenStudy (anonymous):

\[3\sqrt{10}+7\sqrt{15}-6\sqrt{10}-4\sqrt{15}\] is that it?

OpenStudy (anonymous):

well first combine like terms so you get: \[-3\sqrt{10}+3\sqrt{15}\] which is \[3\sqrt{15}-3\sqrt{10}\] thats the second choice ANSWER: SECOND CHOICE